OpenStudy (anonymous):

Take 9 from 6, 10 from 9, 50 from 40 and leave 6. How Come ??

OpenStudy (anonymous):

@mayankdevnani @Fifciol @FibonacciChick666 @Ryaan @betterwiththelightsoff99 @countonme123 @thomaster @satellite73

OpenStudy (anonymous):

@lalaly @uri @phi

OpenStudy (anonymous):

@Loser66 @lissa_bug

OpenStudy (anonymous):

This is pretty vague...

OpenStudy (anonymous):

@genius12 @goldphenoix

OpenStudy (anonymous):

Workhard+playhard=Try hard and solve it yourself..

OpenStudy (anonymous):

@WorkHardPlayHard69 can you tell me any more about the problem?

OpenStudy (anonymous):

@Machida @Mashy @RH @jkbo @gabigurl @emilyhaddad @music101 @MeganEdward

OpenStudy (anonymous):

@WorkHardPlayHard69 You need to explain this better. We can't help you.

OpenStudy (fifciol):

do you mean: \[\frac{ 6 }{ 9 }+\frac{ 9 }{ 10 }+\frac{40 }{ 50}-6\] ?

OpenStudy (anonymous):

@Fifciol yes

OpenStudy (fifciol):

:)

OpenStudy (anonymous):

@dan815 @sexybadass22 @satellite73 @shamim @SA5UK3 @betterwiththelightsoff99 @sexybadass22 @sebasss1234

OpenStudy (anonymous):

@Fifciol @Fifciol @Fifciol @FibonacciChick666 @satellite73 @shamim @shkrina

OpenStudy (anonymous):

@julian25 @jkbo @JOshua_Lewis

OpenStudy (anonymous):

DUDE. Seriously? Stop tagging people. They will see it when they see it.

OpenStudy (anonymous):

@lissa_bug @lissa_bug @lissa_bug

OpenStudy (anonymous):

@satellite73 @shamim @shkrina

OpenStudy (fifciol):

it's -119/30 btw, stop tagging people